# A unified weighting framework for evaluating nearest neighbour classification

### v2

@@ 5.5 Yager-$\frac{1}{2}$ weights
Finally, we consider the results of the Yager-$\frac{1}{2}$ kernel that we have proposed. When we equip NN with both Yager-$\frac{1}{2}$ distance- and rank-weights, this performs slightly better on our data than Samworth distance- and rank-weights, but the difference is not significant ($p<0.25$ across scaling measures). Interestingly, unlike the Samworth kernel, the Yager-$\frac{1}{2}$ kernel appears to perform about as well when only used for distance-weights and when used for both distance- and rank-weights. Correspondingly, Yager-$\frac{1}{2}$ distance-weights perform significantly better than Samworth distance-weights ($p<0.0053$). Thus the main advantage of the Yager-$\frac{1}{2}$ kernel is that it enables comparable performance as the Samworth kernel, but is easier to implement, because it does not require the addition of rank-weights and because it is not dependent on the dimensionality of the dataset.
@@ 6 Conclusion
    Boscovich distance clearly outperforms Euclidean and Chebyshev distance, regardless of other hyperparameter choices.
@@
We hope that any future evaluations of competing proposals will be based on a similar collection of real-life classification problems to the one that we have assembled, which we have made available for this purpose.

### v3

@@ 5.6 Yager-$\frac{1}{2}$ weights
We now consider the results of the Yager-$\frac{1}{2}$ kernel that we have proposed. When we equip NN with both Yager-$\frac{1}{2}$ distance- and rank-weights, this performs slightly better on our data than Samworth distance- and rank-weights, but the difference is not significant ($p<0.25$ across scaling measures). Interestingly, unlike the Samworth kernel, the Yager-$\frac{1}{2}$ kernel appears to perform about as well when only used for distance-weights and when used for both distance- and rank-weights. Correspondingly, Yager-$\frac{1}{2}$ distance-weights perform significantly better than Samworth distance-weights ($p<0.0053$). Thus the main advantage of the Yager-$\frac{1}{2}$ kernel is that it enables comparable performance as the Samworth kernel, but is easier to implement, because it does not require the addition of rank-weights and because it is not dependent on the dimensionality of the dataset.
@@
5.7 Weight optimisation
Finally, we ask whether it is beneficial to optimise not just the number of neighbours $k$ through leave-one-out validation, but also the weight type. Because we obtained our best results with Samworth and Yager weights, we focus on these two weight types for this experiment. In addition, we restrict ourselves to Boscovich distance and $r_{2}$-scaling.
Recall that the Yager kernel is parametrised by $p$, which controls the relative steepness of the function for values close to 0 and close to 1 (Fig. 4). Thus, we can optimise the weight type by optimising $p$. Similarly, the Samworth kernel is parametrised by the value $m$. Until now, we have used the dimensionality of the data to set $m$, but we can also treat it as an independent, optimisable, parameter.
In order to aid the parameter search, we reparametrise the Yager and Samworth kernels by substituting $p=2^{q}$ and $m=2^{-(q+1)}$, resulting, respectively, in $a\mapsto(1-a^{2^{q}})^{2^{-q}}$ and $a\mapsto 1-a^{2^{q}}$. Because we now optimise two hyperparameters, and because $q$ is continuous, unlike $k$, we can no longer perform an exhaustive search. Therefore, we perform Malherbe-Powell optimisation [61, 62, 63] as implemented in the dlib library [64], which was the best-performing algorithm for a series of similar hyperparameter optimisation problems [65], with a search budget of 100 evaluations.
We find that the effectiveness of optimising $q$ depends on a proper delineation of the search space. We originally considered values in $[-10,10]$, however, for the Yager weights most optimised values end up in the relatively narrow interval $[-4,2]$, while for the Samworth weights, there is a large concentration of values at -10, suggesting that their true optimal value is lower. With this search space, optimised Samworth and Yager weights do not outperform their unoptimised equivalents for NN.
However, when we change the search space to $[-4,2]$ for Yager weights and $[-20,0]$ for Samworth weights, we obtain better results. Now, both for a combination of Samworth distance and rank weights ($p=0.045$) as well as for Yager distance weights ($p=0.058$), optimisation increases performance for NN. For FRNN, optimising Samworth distance- and rank weights also increases performance, but the difference is not significant ($p=0.22$). However, FRNN still outperforms NN with optimised Samworth distance and rank weights ($p=0.038$).
@@ 6 Conclusion
    For NN in particular, performance can be further increased by optimising not just $k$, but also the weight type, for which we can use the parameter $p$ of the Yager kernel and the parameter $m$ of the Samworth kernel.
@@
    Boscovich distance clearly outperforms Euclidean and Chebyshev distance, regardless of other hyperparameter choices. Chebyshev distance is a particularly bad choice for very high-dimensional datasets.
  • 7.
@@
We hope that any future evaluations of competing proposals on tabular data will be based on a similar collection of real-life classification problems to the one that we have assembled, which we have made available for this purpose. Finally, it would also be interesting to investigate whether the same weighting kernels perform well for nearest neighbour classification of non-tabular data, like images and texts.
@@ Acknowledgments
Appendix A Proofs that existing weight types can be represented through kernels
In this appendix we will show how most existing weighting proposals discussed in Subsection 2.1 can be rewritten as kernel functions applied to $i^{*}=\frac{i}{k+1}$ (rank-weighting) or $d_{i}^{*}=\frac{d_{i}}{d_{k}}$ (distance-weighting), making them compatible with our generalised definition of NN classification (6). Firstly, note that both in the original equation for NN (1), and in our generalised equation (6), we rescale each class score by the total sum of the weights. Therefore, we do not require that weights sum to 1. Moreover, multiplying all weights by a positive constant produces identical class scores. We will use the proportionality symbol $\propto$ to indicate that two sets of weights only differ by a positive constant factor.
A.1 Linear weights
The linear distance-weights proposed by Dudani [28] can be simplified as follows:
$s_{i}$ $=\frac{d_{k}-d_{i}}{d_{k}-d_{1}}$
$\propto\frac{d_{k}-d_{i}}{d_{k}-d_{1}}\cdot\frac{d_{k}-d_{1}}{d_{k}}$
$=1-d_{i}^{*}.$
Similarly, for the linear rank-weights proposed by Dudani [28], we have:
$w_{i}$ $=k-i+1$
$\propto\frac{k-i+1}{k+1}$
$=1-i^{*}.$
And likewise for the linear rank-weights proposed by Stone [30]:
$w_{i}$ $=\frac{k-i+1}{k(k+1)/2}$
$\propto\frac{k-i+1}{k+1}$
$=1-i^{*}.$
A.2 Sugeno weights
In a similar way, we can simplify the weights proposed by Gou et al. [23]:
$s_{i}$ $=\frac{d_{k}-d_{i}}{d_{k}-d_{1}}\cdot\frac{d_{k}+d_{1}}{d_{k}+d_{i}}$ (7)
$=\frac{1-d_{i}^{*}}{1+d_{i}^{*}}.$
We call this the Sugeno-1 kernel, because it is in fact an instance of fuzzy Sugeno negation [66], with $\lambda=1$:
$a\longmapsto\frac{1-a}{1+\lambda a}$
A.3 Quadratic weights
For the quadratic rank-weights proposed by Stone [30], we are left with a remainder, but this vanishes as $k$ increases:
$w_{i}$ $=\frac{k^{2}-(i-1)^{2}}{k(k+1)(4k-1)/6}$
$\propto\frac{k^{2}-(i-1)^{2}}{(k+1)^{2}}$
$=1-\frac{2k+i^{2}-2i+2}{(k+1)^{2}}$
$=1-(i^{*})^{2}-\frac{2(k-i+1)}{(k+1)^{2}}$
$\rightarrow 1-(i^{*})^{2}\text{ as }k\rightarrow\infty.$
A.4 Reciprocally linear weights
The reciprocally linear distance-weights proposed by Dudani [28] are proportional to a reciprocally linear kernel:
$s_{i}$ $=\frac{1}{d_{i}}$
$\propto\frac{d_{k}}{d_{i}}$
$=\frac{1}{d_{i}^{*}}.$
Note that this kernel is improper, as its value goes to $\infty$ as $d_{i}^{*}$ goes to 0.
Likewise, for the reciprocally linear rank-weights proposed by Gou et al. [32], we have:
$w_{i}$ $=\frac{1}{i}$
$\propto\frac{k+1}{i}$
$=\frac{1}{i^{*}}.$
A.5 Reciprocally quadratic weights
Analogously, the reciprocally quadratic distance-weights proposed by Shepard [26] are proportional to a reciprocally quadratic kernel, which is also improper:
$s_{i}$ $=\frac{1}{d_{i}^{2}}$
$\propto\frac{d_{k}^{2}}{d_{i}^{2}}$
$=\frac{1}{\left(d_{i}^{*}\right)^{2}}.$
A.6 Macleod weights
We can also simplify the distance-weights proposed by Macleod et al. [37]:
$s_{i}$ $=\frac{d_{k}-d_{i}+d_{k}-d_{1}}{2(d_{k}-d_{1})}$
$=2-d_{i}^{*}-d_{1}^{*}.$
A.7 Laplace weights
A.8 Samworth weights
For the Samworth weights, if we fix a particular value $k>0$, and write $h=\frac{1}{k+1}$ (for reasons of space), we can define the following kernel function $f^{k}$ such that $f^{k}(\frac{i}{k+1})$ is the $i$th weight as in (2):
Lemma 1.
$\lim_{k\rightarrow\infty}\frac{2k}{m+2}\cdot f^{k}=1-a^{\frac{2}{m}}.$
Proof.
$\mathrel{\phantom{=}}\lim_{k\rightarrow\infty}\frac{2k}{m+2}\cdot f_{k}$
$=1-\frac{m}{m+2}\cdot\left(1+\frac{2}{m}\right)a^{\frac{2}{m}}$ (*)
$=1-a^{\frac{2}{m}},$
where ()(*)( ∗ ) is the polynomial rule for derivation. ∎
Accordingly, we call $1-a^{\frac{2}{m}}$ the Samworth kernel. Note that the Samworth kernel is a fuzzy negation, since it sends $a=1$ to 0.
A.9 Fibonacci weights
The only rank-weights that cannot be obtained by applying a kernel function are the Fibonacci weights from Pao et al. [33], because their relative distribution depends on $k$.